Session: Feb 2003 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Fiocchi at Tue Feb 25 18:51:29 2003) Success :
Good weather, clear sky until the end.

Jodrell_Bank (Z.P. (for Alastair Gunn) at Thu Feb 27 08:59:53 2003) Success :
Joined in the observations in the last minute for the request of JIVE. Recorded on tape only. The telescope was off source after 17:00 UT. There were no apparent problems during the experiment apart from the usual L-band RFI.

Westerbork (Z.P. (for Tony Foley)) at Thu Feb 27 09:09:52 2003) Minor failures : LATE
Started late by about a minute. Severe RFI during the first scan.

Noto (P. Cassaro at Thu Feb 27 09:41:58 2003) Minor failures : TSYS
From 14:36 to 15:15 UT the tsys has not been calculated due to errors (now fixed) in the midob procedure